OpenStudy (anonymous):

Solve the system. y = -1/3 x + 2 and x + 3y = 3 (0, 1) (1, 0) (3, 1/3) (3/2, 0) no solution The ordered pair (-1, -1) is a solution of the system -4x + 2y = 2 and x + y = -2. (Points : 1) True False How many solutions exist for the system y = 2x + 5 and y = -0.2x - 4? (Points : 1) no solution one solution infinitely many solutions How many solutions exist for the system, x + y = 1 and y = -x + 1? (Points : 1) no solution one solution infinitely many solutions

OpenStudy (anonymous):

Last questions please thank you!

OpenStudy (lgbasallote):

FIRST QUESTION: y = -1/3 x + 2 and x + 3y = 3 multiply the first equation by 3 3y = -x +2 so x+3y = 2 x+3y = 3 no solution (this is parallel) SECOND QUESTION: substitute the ordered pair to both equations...if the coordinates are correct when substituted, then they are the ones. THIRD QUESTION: Use discriminant(b^2 -4ac) FOURTH QUESTION: same with third